OpenStudy (anonymous):

Find the equation of a line given the point and slope below. Arrange your answer in the form y = mx + b, where b is the constant. (-6, 1) m = 6 i got y=37?

OpenStudy (anonymous):

b=37

OpenStudy (anonymous):

can you learn from your previous questions? y = mx + b 1 = -36 + b b = 37 y = 6x + 37

OpenStudy (anonymous):

y-1=6(x+6) y=6x+36+1=6x+37